your neighbor leaves his yard, which is 120 ft from the bus stop, and begins to walk toward you at the bus stop, but passes you. He is walking at a speed of 4ft/s. his distance d from you after t seconds is given by d=|120-4t|. after how many seconds is your neighbor 30 ft from you?
1
Expert's answer
2020-12-17T07:15:30-0500

d = (120-4t)

The neighbour can be 30ft away from me before he passes me and also 30ft away from me after he passes me.

So, d can be 30 or -30

30 = 120-4t

4t = 120-30

4t = 90

t = 90/4 = 22.5s

Also,

-30 = 120-4t

4t = 150

t = 150/4 = 37.5s